OpenStudy (anonymous):

1/5 (p - 5) = 3/5p + 1/10p + 1 I got two different answers each time I did it. 1.) p = 101/13 & 2.) p = -11/5 Show steps, please? :)

OpenStudy (anonymous):

I agree with your second answer - Here's what I did: First distribute: \[\frac{ 1 }{5}p -1 = \frac{ 3 }{ 5 }p+\frac{ 1 }{ 10}p +1\] Then I could either get a common denomenator for all the fractions (change them into 10ths), or just multiply both sides by 10 to get rid of fractions. 2p - 10 = 6p + p + 1 combine like terms to get: -11 = 5p \[p=\frac{ -11 }{ 5 }\]